OpenStudy (anonymous):

Find the exact value of sin-1 (-square root 3/2) Write your answer in radians in terms of pi.

OpenStudy (lgbasallote):

this is a trig identity

OpenStudy (anonymous):

okay so whats the answer? please?

OpenStudy (lgbasallote):

 sin is opposite over hypotenuse right? from my drawing...which angle has opposite of sqrt 3?

OpenStudy (lgbasallote):

this is just easy...just look at the drawing

OpenStudy (lgbasallote):

which angle is opposite \(\sqrt 3\)

OpenStudy (anonymous):

so whats the answer?

OpenStudy (lgbasallote):

-___- the angle opposite sqrt 3

OpenStudy (lgbasallote):

it's just some sense of direction..dont tell me you need guidance for that

OpenStudy (anonymous):

just tell me the answer

OpenStudy (anonymous):

-pi/3 (-60 degrees) Solution: \[sin(-pi/3)=-sin(pi/3)\]\[=-\sqrt{3}/2\] \[arcsin-\sqrt{3}/2 = -pi/3\]

OpenStudy (anonymous):

okay thanks!